# Geometric dependence of exchange bias in tilted three-dimensional CoFe/IrMn microwires

### v1

@@ I Introduction
In this work, we investigate EB in 3D microwires with varying inclinations, focusing on how the directional nature of sputter deposition affects the effective FM and AF thicknesses, as well as the interfacial properties of CoFe/IrMn bilayers. By comparing smooth and rougher 3D scaffolds, together with planar bilayers of systematically varied layer thicknesses, we assess the roles of geometric thickness variations, surface roughness and microstructural effects in determining the exchange-bias field and coercivity in nonplanar architectures.
@@ II.2 Effect of tilted 3D geometry on exchange bias
Oblique-angle deposition is known to modify the film microstructure, promoting columnar growth and increased interface roughness [2]. Fleischmann et al. [19] demonstrated that enhanced interface roughness in exchange-biased FM/AF systems increases coercivity by introducing additional pinning sites, while the exchange-bias field exhibits a nonmonotonic dependence on roughness due to competing effects of uncompensated spins and interfacial disorder. As discussed in the following section, in our case, we understand the increase in $H_{\mathrm{EB}}$ for inclined wires to be primarily attributed to a reduction in FM thickness, indicating that microstructural changes have a minor effect on $H_{\mathrm{EB}}$. Nevertheless, the more complex magnetization reversal observed for the 45° wire suggests enhanced domain-wall pinning, which may arise from either the modified film microstructure or the reduced FM thickness locally at the wire edges.
@@ II.4 Conclusions
This work demonstrates the realization of exchange-biased 3D CoFe/IrMn microwires by integrating TPL-fabricated microwire scaffolds of inclination angles up to $45^{\circ}$ with magnetron sputtered CoFe/IrMn bilayer. DF-MOKE measurements reveal a systematic increase of $H_{\mathrm{EB}}$ and $H_{\mathrm{c}}$ with wire inclination, which is primarily attributed to the geometry-induced reduction of the FM layer thickness, a result of the directional deposition on a 3D tilted geometry. Analysis of surface morphology and the variation of $H_{\mathrm{EB}}$ with both geometric FM thicknesses ($t^{\mathrm{FM}}_{G}$) and nominal FM thicknesses ($t^{\mathrm{FM}}_{0}$) indicates that the interface quality of FM/AF bilayer deposited on smooth 3D wire scaffolds remains largely unaffected by the 3D geometry or the film microstructure, highlighting the compatibility of 3D geometries with multilayered stacks with functional interfaces grown by physical vapour deposition methods.

### v2

@@ I Introduction
In this work, we investigate EB in 3D microwires with varying inclinations, focusing on how the directional nature of sputter deposition affects the effective FM and AF thicknesses, as well as the interfacial properties of CoFe/IrMn bilayers. By comparing smooth and rougher 3D scaffolds, together with planar bilayers of systematically varied layer thicknesses, we assess the roles of geometric thickness variations, surface roughness and microstructural effects in determining the EB field and coercivity in nonplanar architectures.
@@ II.1 Exchange-biased 3D CoFe/IrMn microwires
Figure 2: Fabrication of exchange-biased 3D CoFe/IrMn microwires. (a) Schematic illustration of the TPL fabrication of 3D microwire scaffolds on a glass substrate, polymerization occurring at the focus point of the laser within the photoresist. (b) Schematic of the magnetron sputter deposition performed in the presence of an in-plane magnetic field applied along $x$-axis (see Fig. 1a for the coordinate systems). (c) Illustration of the resulting possible CoFe magnetization orientation in planar film and the 3D wire as a result of deposition in the presence of a magnetic field, followed by the IrMn deposition. The inset shows schematically the possible spin orientations of the CoFe/IrMn bilayer. (d-f) SEM images of the 3D microwires with inclinations of 0°, 30° and 45° (with respect to the substrate) after the deposition of the stack.
@@ II.2 Effect of tilted 3D geometry on exchange bias
Oblique-angle deposition is known to modify the film microstructure, promoting columnar growth and increased interface roughness [2]. Fleischmann et al. [19] demonstrated that enhanced interface roughness in exchange-biased FM/AF systems increases coercivity by introducing additional pinning sites, while the $H_{EB}$ exhibits a nonmonotonic dependence on roughness due to competing effects of uncompensated spins and interfacial disorder. As discussed in the following section, in our case, we understand the increase in $H_{\mathrm{EB}}$ for inclined wires to be primarily attributed to a reduction in FM thickness, indicating that microstructural changes have a minor effect on $H_{\mathrm{EB}}$. Nevertheless, the more complex magnetization reversal observed for the 45° wire suggests enhanced domain-wall pinning, which may arise from either the modified film microstructure or the reduced FM thickness locally at the wire edges.
@@ II.3 Role of surface roughness and tilted geometry on bilayer growth and magnetic performance
Figure 4: Exchange bias and coercivity as a function of wire geometry and surface roughness. (a) EB and (b) Coercivity as a function of inverse geometric FM thickness ($1/t_{G}^{\mathrm{FM}}$) for FM- and AF-series. Solid symbols correspond to planar film data, while open symbols represent wire data. Dashed lines indicate linear fits to the wire data at the three inclination angles, whereas the solid black line shows the linear fit to the planar film data in the FM series, with the dotted black line representing its extrapolation to lower $t_{0}^{\mathrm{FM}}$ values. $\sigma_{\mathrm{Planar\ films/Wires}}^{\mathrm{CoFe}\mathit{X}/\mathrm{IrMn}\mathit{Y}}$ (in $\mathrm{mT\,nm}$) denotes the slope obtained from the linear fit to the “Planar films/Wires” data points of the $\mathrm{CoFe}\mathit{X}/\mathrm{IrMn}\mathit{Y}$ bilayer. The labels $0^{\circ}$, $30^{\circ}$ and $45^{\circ}$ associated with the wire data points indicate the corresponding wire inclination angles. (c) SEM images of 0° wires (first column) for (1) CoFe5.5/IrMn14, (2) CoFe7/IrMn14 and (3) CoFe5.5/IrMn5, with corresponding AFM topographies (second column) acquired at indicated marker positions. The height color scale is shared across AFM images. The third column shows magnified views of 45° wires. (d) Power spectral density (PSD) vs. spatial frequency $k$, with planar films shown as solid lines and wires as dashed lines. (e) Slopes, extracted from linear fits of $H_{\mathrm{EB}}$ versus $1/t_{G}^{\mathrm{FM}}$, are plotted as a function of RMS roughness. Roughness data correspond to the planar geometries: the $0^{\circ}$ wire and the average of the $\mathrm{CoFe}\mathit{X}/\mathrm{IrMn}14$ planar films, with the standard deviation from this averaging used as the error bar. Marker symbols and colors, as indicated in the legend at the top, are consistent across all panels.
@@ II.4 Conclusions
This work demonstrates the realization of exchange-biased 3D CoFe/IrMn microwires by integrating TPL-fabricated microwire scaffolds of inclination angles up to $45^{\circ}$ with magnetron sputtered CoFe/IrMn bilayer. DF-MOKE measurements reveal a systematic increase of $H_{\mathrm{EB}}$ and $H_{\mathrm{c}}$ with wire inclination, which is primarily attributed to the geometry-induced reduction of the FM layer thickness, a result of the directional deposition on a 3D tilted geometry. Analysis of surface morphology and the variation of $H_{\mathrm{EB}}$ with both geometric FM thicknesses ($t^{\mathrm{FM}}_{G}$) and nominal FM thicknesses ($t^{\mathrm{FM}}_{0}$) indicates that the interface quality of FM/AF bilayer deposited on smooth 3D wire scaffolds remains largely unaffected by the 3D geometry or the film microstructure, highlighting the compatibility of 3D geometries with multilayered stacks with functional interfaces grown by physical vapor deposition methods.
@@ Acknowledgments
Supporting Information
Geometric dependence of exchange bias in tilted three-dimensional CoFe/IrMn microwires
Balram Singh1, Aman Singh2,3, Stefan Mikulik1, Jakub Jurczyk1, Volker Neu2, and Amalio Fernández-Pacheco1
1 Institute of Applied Physics, TU Wien, 1040 Vienna, Austria
2 Leibniz Institute for Solid State and Materials Research Dresden, 01069 Dresden, Germany
3 Institute of Applied Physics, TU Dresden, 01187 Dresden, Germany
Fig. S1. Magnetization reversal in FM/AF bilayer systems. DF-MOKE hysteresis loops for (A) CoFe7/IrMn14, (B) CoFe5.5/IrMn5, (C) CoFe8.5/IrMn14 and (D) CoFe5.5/IrMn3. The magnetic field ($\mu_{0}H_{x^{\prime}}$) is applied along the $x^{\prime}$ direction (see Fig. 1a, for the coordinate systems).
Fig. S2. AFM and SEM imaging of FM/AF systems. AFM topography images of FM/AF planar films and SEM images of 3D FM/AF wires are shown for (A) CoFe5.5/IrMn14, (B) CoFe7/IrMn14, (C) CoFe5.5/IrMn5 and (D) CoFe8.5/IrMn14.